INFRARED SPECTROSCOPY OF SIZE-SELECTED CARBOCATIONS II.: C_{6}H_{7}^{+}, C_{7}H_{7}^{+}, AND C_{7}H_{9}^{+}

\underline{GARY E. DOUBERLY}, ALLEN M. RICKS, and MICHAEL A. DUNCAN, Department of Chemistry, University of Georgia, Athens, Georgia 30602.

Cold, rare gas tagged, gas phase carbocations are produced in a pulsed electric discharge supersonic expansion cluster source. The infrared spectra of the size-selected species are obtained via infrared photodissociation spectroscopy. The structure and spectroscopy (700-4000 cm^{−1}) of the benzenium, toluenium and tropylium ions will be discussed. The infrared spectral signatures of the title ions are compared to the interstellar unidentified infrared emission bands.